OpenStudy (ny,ny):

What is the solution of the equation 2log_4(5x) = 3?

OpenStudy (ny,ny):

@ganeshie8

OpenStudy (jojokiw3):

Ah okay. So, log rules. \[a \log b = \log b^{a}\] and \[(\log_{4} a = b) = (a = 4^b)\] Can you try to figure it out with these?

OpenStudy (jojokiw3):

You want to try and isolate 5x :)

OpenStudy (ny,ny):

Okay. 5x^2 = 4^3 ?

OpenStudy (jojokiw3):

Yes. Keep going.

OpenStudy (ny,ny):

\[x = \sqrt{64/5}\]

OpenStudy (jojokiw3):

Hm. No. Ah I see what you did: \[(5x)^2 = 25x^2\] first.

OpenStudy (jojokiw3):

Do you have the answer? @NY,NY

OpenStudy (jojokiw3):

Haha you have 1 tiny mistake.

OpenStudy (ny,ny):

x = 8/5 *

OpenStudy (jojokiw3):

Yep! Good job!

OpenStudy (ny,ny):

Thank you :) @jojokiw3

OpenStudy (jojokiw3):

No problem. ^-^